OpenStudy (anonymous):

please :) Choose the correct simplification of the expression a2• a3

OpenStudy (anonymous):

a^5

OpenStudy (diyadiya):

\[a^2.a^3= a^{2+3}=a^5\]

OpenStudy (apoorvk):

\[x^a * x^b = x^{a+b}\]

OpenStudy (apoorvk):

hence here, you should get a^(2+3) = a^5